OpenStudy (anonymous):

find all the solutions to the equation in the intercal [0,2pi) using the equation: sec^2x+5tanx=-2

OpenStudy (anonymous):

sec^2x = 1 + tan^2x so tan^2x + 5 tanx + 3 = 0 solving this give tan x = -0.6972 or -4.302 values for x are 2.53, 5.67, 1.8, 4.94 radians if you require answers is degrees multiply the above by 180/pi

OpenStudy (anonymous):

ok so i can get to the tan^2 + 5tanx = 0 but how do you actually get to the answers?

OpenStudy (anonymous):

you can solve that quadratic using the formula. find the angles by plugging tan-1 0.6972 and 4.302 into your calculator this gives angles in first quadrant - then work out angles corresponding to negatives - tangents are negative in the second and fourth quadrants

OpenStudy (anonymous):

isnt ther something that you can do with it by factoring the left side of tan^2x+5tanx+3=0? and then you set them equal to 0 and solve?

OpenStudy (anonymous):

hey I have one more question how do u get the values of x?